# Global Prediction of COVID-19 Variant Emergence Using Dynamics-Informed Graph Neural Networks

### v1

@@ IV-B Benchmark Pipeline
In the context of emergence prediction, our objective is to retrospectively predict whether a given variant will become dominant in each country. Given the presence of class imbalance in our data, we employ the Macro F1-score as an evaluation metric, which considers both precision and recall across all classes, providing a balanced assessment [47, 48].
@@ V-A Results Analysis
The observations from Figure 5(a) give us two key insights. Firstly, the primary source of error is associated with variants that persist for an extended duration, such as 21J.Delta or 23E.Omicron. Examining the color gradients, these variants appear to linger in only a few countries before spreading to others. In reality, variants typically do not endure for such prolonged periods without either being superseded by another variant or transmitting to additional countries [38, 49]. This suggests a potential issue in the data capturing these lingering variants. Secondly, despite this being a major source of error for GCN-based models, they exhibit overall robustness to outlier predictions, unlike the dynamics-based model as seen in Figure 5(b). Also, note that our ground truth is at a granularity of 14 days which may inflate the error. Suppose, in reality, the first day of a variant reaching a desired prevalence threshold is $t=14k+1$, for some integer $k$. In the ground truth, it will appear at the next multiple of 14, i.e., $t+13$. If our prediction was $t^{\prime}=14k=t-1$, the actual error is 1 day, while in comparison to the observed ground truth, it will result in an error of 14 days.
@@
While the dynamics-based model does output very good predictions, it fails often and dramatically in some cases. Figure 7 presents examples of 3D points in space utilized for fitting linear models. In Figure 6(a), the 3D points form a space where a plane can be fitted exceptionally well. However, as newer COVID variants emerge, the dynamics may evolve, causing some initial assumptions, such as full cross-immunity [38], to become invalid. This evolution makes it more challenging to perfectly fit a linear model, as depicted in Figure 6(c).

### v2

@@ IV-B Benchmark Pipeline
In the context of emergence prediction, our objective is to retrospectively predict whether a given variant will become dominant in each country. Given the presence of class imbalance in our data, we employ the Macro F1-score as an evaluation metric, which considers both precision and recall across all classes, providing a balanced assessment [opitz2019macro, lipton2014optimal].
@@ V-A Results Analysis
The observations from Figure 6(a) give us two key insights. Firstly, the primary source of error is associated with variants that persist for an extended duration, such as 21J.Delta or 23E.Omicron. Examining the color gradients, these variants appear to linger in only a few countries before spreading to others. In reality, variants typically do not endure for such prolonged periods without either being superseded by another variant or transmitting to additional countries [Chen2023, Verywell_Health_2023]. This suggests a potential issue in the data capturing these lingering variants. Secondly, despite this being a major source of error for GCN-based models, they exhibit overall robustness to outlier predictions, unlike the dynamics-based model as seen in Figure 6(b). Also, note that our ground truth is at a granularity of 14 days which may inflate the error. Suppose, in reality, the first day of a variant reaching a desired prevalence threshold is $t=14k+1$, for some integer $k$. In the ground truth, it will appear at the next multiple of 14, i.e., $t+13$. If our prediction was $t^{\prime}=14k=t-1$, the actual error is 1 day, while in comparison to the observed ground truth, it will result in an error of 14 days.
@@
While the dynamics-based model does output very good predictions, it fails often and dramatically in some cases. Figure 7 presents examples of 3D points in space utilized for fitting linear models. In Figure 7(a), the 3D points form a space where a plane can be fitted exceptionally well. However, as newer COVID variants emerge, the dynamics may evolve, causing some initial assumptions, such as full cross-immunity [Chen2023], to become invalid. This evolution makes it more challenging to perfectly fit a linear model, as depicted in Figure 7(c).